remi2019:

Question 17 Maria races her bicycle for 94m. A wheel of her bicycle turns 47 times as the bicycle travels this distance. What is the diameter of the wheel? Use the value 3.14 for pie. Round your answer to the nearest tenth. Do not round any intermediate steps.

snowflake0531:

this means that the wheel circumference multiplied by 47 equals 94 so, (dpi)47 = 94 do you know how to simplify it to get d?

snowflake0531:

anyways, you first divide 47 be each side, resulting dpi=2 bc pi is 3.14 you divide 2 by 3.14 since it wants you to round to the nearest tenths you can round 0.63694267515 to 0.6 hm @AZ, i feel like i did to wrong lol

snowflake0531:

*it

AZ:

yeah, that's correct we know that the circumference of her bike times how many times it turned is the total distance she travelled so circumference * 47 = 94 and the circumference formula is C = 2* \(\pi\) * radius or you could remember that 2 * radius is the diameter and get C = \(\pi\) * diameter

AZ:

so you're on the right track :)

remi2019:

2*47=94

AZ:

so that means the circumference is 2 and so we have circumference = 2 * pi and we know pi is a constant which rounds to 3.14

remi2019:

6.28

AZ:

oops sorry

AZ:

I made a mistake

AZ:

circumference is 2 2 = pi * diameter

remi2019:

0.636942675159

snowflake0531:

then round to the nearest tenth

remi2019:

0.6

snowflake0531:

ig lol